OpenStudy (anonymous):

Solve: Square root x=4

OpenStudy (anonymous):

\[\sqrt{x}=4\]

OpenStudy (anonymous):

\[\sqrt{x}=4\] \[\sqrt{x}^{2}=4^2\] x=16